OpenStudy (anonymous):

5. A spinner is divided into four equal sections--red, blue, orange and green. You spin twice. What is the probability you will spin blue and then orange? The probability is . 6. A spinner is divided into four equal sections--red, blue, orange and green. What is the probability you will spin blue or orange on your first spin? The probability is

OpenStudy (lena772):

1/4*1/4=1/16

OpenStudy (lena772):

^^ for 5

OpenStudy (lena772):

2/4=1/2 (for 6)

OpenStudy (abhisar):

You probably didn't notice, but you posted your \(\large\sf\color{blue}{\underline{\href{ /study#/groups/mathematics}{Mathematics}}}\) question in the Biology section. Next time when you ask a question, check if you're in the right section :)